OpenStudy (anonymous):

Madeline hosted a dinner party at her favorite restaurant. She had to pay $15 per guest. The restaurant took $60 off the bill due to a mistake in the order. If the final bill is $675, how many guests attended the dinner party? 39 41 45 49

OpenStudy (anonymous):

60+675=735/15=49. 60 got subtracted from initial value, they gave final value, hence the 60+675.